OpenStudy (anonymous):

Can someone help me with this? Find the roots of the following problem, x^2+7x+12

OpenStudy (anonymous):

*following quadratic by factoring

OpenStudy (solomonzelman):

\(\large\color{black}{ \displaystyle x^2+7x+12=0}\) \(\large\color{black}{ \displaystyle x^2+3x+4x+12=0}\) \(\large\color{black}{ \displaystyle x(x+3)+4(x+3)=0}\) and so forth...